OpenStudy (anonymous):

5) Find the x- and y-intercept of the line. 5x+10y=30

OpenStudy (anonymous):

@aliboo997

OpenStudy (eric_d):



OpenStudy (eric_d):

Arrange it in that form

OpenStudy (igreen):

The y-intercept is what y is when x = 0, so plug in x = 0 in the equation: 5x + 10y = 30 5(0) + 10y = 30 0 + 10y = 30 10y = 30 y = 3

OpenStudy (igreen):

The x-intercept is what x is when y = 0, so plug in y = 0 in the equation: 5x + 10y = 30 5x + 10(0) = 30 5x + 0 = 30 5x = 30 x = 6 So the y-intercept is (0, 3) and the x-intercept is (6, 0).